The bases of an isosceles trapezoid are 12 inches and 18 inches, and the legs are 5

inches. A rectangle equal in area to the trapezoid has one dimension 6 inches.

Find the other dimension.



1
Expert's answer
2021-02-02T06:20:29-0500

Find the height of the isosceles trapezoid "h" from the Pythagorean Theorem


"5^2=h^2+(\dfrac{18-12}{2})^2"

Then "h=4\ in"

The area of the trapezoid


"A_t=(\dfrac{18+12}{2})h=15(4)=60\ {in}^2"

The area of the rectangle


"A_r=length\times width=6\times width"

Given "A_t=A_r"


"6\times width=60"

"width=10\ in"

A rectangle is equal in area to the trapezoid has another dimension of 10 inches.